# Understanding Optimal Feature Transfer via a Fine-Grained Bias-Variance Analysis

### v1

@@ Abstract
In the transfer learning paradigm models learn useful representations (or features) during a data-rich pretraining stage, and then use the pretrained representation to improve model performance on data-scarce downstream tasks. In this work, we explore transfer learning with the goal of optimizing downstream performance. We introduce a simple linear model that takes as input an arbitrary pretrained feature transform. We derive exact asymptotics of the downstream risk and its fine-grained bias-variance decomposition. Our finding suggests that using the ground-truth featurization can result in “double-divergence” of the asymptotic risk, indicating that it is not necessarily optimal for downstream performance. We then identify the optimal pretrained representation by minimizing the asymptotic downstream risk averaged over an ensemble of downstream tasks. Our analysis reveals the relative importance of learning the task-relevant features and structures in the data covariates and characterizes how each contributes to controlling the downstream risk from a bias-variance perspective. Moreover, we uncover a phase transition phenomenon where the optimal pretrained representation transitions from hard to soft selection of relevant features and discuss its connection to principal component regression.
@@ 1 Introduction
Despite the widespread adoption of this paradigm, the mechanisms that produce effective representations remain poorly understood. In particular, one may ask the following question: for a given set of downstream tasks, what constitutes optimal pretrained representation? Intuitively, this representation must: (1) acquire features that are broadly useful across downstream tasks, and (2) understand the structure of the data covariates. Here we study these mechanisms theoretically in a simple setting. We find that both mechanisms are significant and we describe their respective roles in controlling the downstream risk from a bias-variance perspective.
@@ 1.1 Contributions
    Introducing a simple, exactly solvable, regression-based model that interpolates from soft/no featurization to strong pretrained featurization by varying regularization parameters.
@@
    Deriving the exact asymptotics for the downstream risk and its fine-grained bias-variance decomposition given an arbitrary linear representation from upstream. We find that using the ground-truth featurization can result in “double-divergence” of the asymptotic risk, indicating that it is not necessarily optimal for downstream performance.
@@
    Proposing optimization methodologies to minimize the asymptotic downstream risk as a function of the pretrained representation; conducting ablation studies comparing the total risk, bias, and variance of the optimally pretrained predictor with predictors with no featurization and ground-truth featurization. We also adapt the methodology for a minimax objective that controls the worst-case performance among downstream tasks.
@@
    Finding that learning both task-relevant features and structures in data covariates are vital, by interpreting the structure of optimal pretraining. Our analysis also reveals a novel phase transition phenomenon in which the optimal pretrained representation transitions from hard to soft selection of relevant features; we discuss this phenomenon in relation to principal component regression.
@@ 2.1 Settings
In this paper, we examine a simple setting where downstream tasks share a common linear representation that can be learned upstream. That is, we assume downstream tasks are supervised learning tasks where the responses are generated i.i.d. from
$y\sim{\mathbf{x}}^{\top}{{\bm{\beta}^{\star}}}+\varepsilon,\qquad{\bm{\beta}^{\star}}=\mathbf{B}^{\star}\bm{\alpha}^{\star}$ (1)
where $\mathbf{x}\sim N(\bm{0},\mathbf{\Sigma})$ are Gaussian covariates with fixed covariance $\mathbf{\Sigma}$ and the unknown parameter ${\bm{\beta}^{\star}}$ characterizes the particular downstream task. The ground-truth representation $\mathbf{B}^{\star}$ is common to all downstream tasks and can be learned upstream. This assumption of shared linear representation follows a long line of past and concurrent work which we review in Section 2.3
For each downstream task, the objective is to predict the label ${y}_{\mathrm{new}}$ of an i.i.d. sample of covariates $\mathbf{x}_{\mathrm{new}}$ with a certain predictor $\hat{y}$ and minimizes the task-specific risk  [^1] Note we assume ${y}_{\mathrm{new}}$ does not have label noise.
$R:=\mathbb{E}_{\mathbf{x}_{\mathrm{new}}}\left({y}_{\mathrm{new}}-\hat{y}\right)^{2}=\mathbb{E}_{\mathbf{x}_{\mathrm{new}}}\left(\mathbf{x}_{\mathrm{new}}^{\top}{\bm{\beta}^{\star}}-\hat{y}\right)^{2}.$
We consider the few-shot setting, where the amount of task-specific downstream data $n$ is scarce. In particular, one may have $n\ll p$. To overcome this deficiency, one can learn the shared linear representation $\mathbf{B}^{\star}$ from upstream data and subsequently incorporate it into the downstream estimation of ${{\bm{\beta}^{\star}}}$. We give an example below where $\mathbf{B}^{\star}$ and $\bm{\alpha}^{\star}$ have clear interpretations.
Example 2.1.
At the pretraining stage, the model receives training data for $q$ distinct upstream tasks $\quantity(\mathbf{y}_{\mathsf{pre}}^{(i)},\mathbf{X}_{\mathsf{pre}}^{(i)})_{i=1}^{q}$ where
@@ 2.2 Ridgeless Regression with Pretrained Representation
In the following, we assume that the model has already acquired a representation matrix $\widehat{\mathbf{B}}\in\mathbb{R}^{p\times k}$ from the upstream pretraining. Definition 2.2 below introduces a downstream predictor using this pretrained representation. In Section 4, we will optimize over this representation matrix $\widehat{\mathbf{B}}$. Before then, $\widehat{\mathbf{B}}$ should be regarded as a fixed, deterministic feature transform.
@@
For a downstream task task with training data $(\mathbf{X},\mathbf{y})$, define the empirical loss
@@
$\hat{y}(\mathbf{x}_{\mathrm{new}})=\mathbf{x}_{\mathrm{new}}^{\top}\hat{\bm{\beta}}\quad\text{where}\quad\quantity(\hat{\bm{\beta}},\hat{\bm{\alpha}})=\lim_{\lambda_{0}\to 0}\operatornamewithlimits{argmin}_{\bm{\beta},\bm{\alpha}}\mathcal{L}(\bm{\beta},\bm{\alpha}).$
Remark 2.3 (Strong and Weak Featurization).
A special feature of the loss function Eq. (2) is that it contains a penalty term $\lambda\|\bm{\beta}-\widehat{\mathbf{B}}\bm{\alpha}\|_{2}^{2}$ that softly enforces the featurization $\hat{\bm{\beta}}=\widehat{\mathbf{B}}\hat{\bm{\alpha}}$. The strength of featurization is controlled by regularization parameters $\lambda_{0},\lambda_{\mathbf{\alpha}},\lambda_{\mathbf{\beta}},\lambda$. One can adjust these parameters to interpolate from strong featurization to weak or no featurization. For instance, we can consider a strong-featurization limit by letting $\lambda_{\mathbf{\beta}}\to 0$ and $\lambda\to+\infty$, we get $\hat{\bm{\Gamma}}^{-1}\to\lambda_{\mathbf{\alpha}}^{-1}\widehat{\mathbf{B}}\widehat{\mathbf{B}}^{\top}$, and it follows from Eq. (5) that
@@
Therefore, in this limit, $\hat{y}$ makes predictions by regressing with $\hat{\bm{\alpha}}_{0}$ composed with the featurization generated by $\widehat{\mathbf{B}}$, which is reminiscent of final-layer head-tuning in neural networks [5, 75, 49]. In fact, $\hat{\bm{\alpha}}_{0}$ is the minimum $\ell_{2}$-norm solution, and by a well-known result (see e.g. [26]), it is also the limit of gradient flow on the objective $\|\mathbf{y}-\mathbf{X}\widehat{\mathbf{B}}\bm{\alpha}\|^{2}_{2}$ when initialized at zero.
Meanwhile, the standard ridgeless predictor with no featurization can be recovered from Definition 2.2 by effectively removing the penalty term $\lambda\|\bm{\beta}-\widehat{\mathbf{B}}\bm{\alpha}\|_{2}^{2}$ linking $\bm{\alpha}$ and $\bm{\beta}$ in $\mathcal{L}$. This is achieved by setting $\lambda\to 0$, which implies $\hat{\bm{\Gamma}}\to\lambda_{\mathbf{\beta}}\mathbf{I}_{p}$, and the predictor is given by Eq. (5). Alternatively, assuming $\hat{d}_{i}>0$ for all $i=1,\ldots,p$ and taking $\lambda_{\mathbf{\alpha}}\to 0$, we again see $\hat{\bm{\Gamma}}\to\lambda_{\mathbf{\beta}}\mathbf{I}_{p}$ since $\hat{\bm{\alpha}}$ is free to exactly solve $\bm{\beta}=\widehat{\mathbf{B}}\bm{\alpha}$ for any $\bm{\beta}$ without constraints on its norm.
Note we have restricted ourselves to the ridgeless limit, $\lambda_{0}\to 0$, in which explicit expressions for $\quantity(\hat{\bm{\beta}},\hat{\bm{\alpha}})$ are possible. To present these expressions, we first write the singular value decomposition of $\widehat{\mathbf{B}}$ as
@@
Remark 2.6 (Connection to Generalized Ridge Regression).
One may rewrite $\hat{\bm{\beta}}$ in Eq. (5) as:
$\hat{\bm{\beta}}=\lim_{\lambda_{0}\to 0}\operatornamewithlimits{argmin}_{\bm{\beta}}\norm{\mathbf{y}-\mathbf{X}\bm{\beta}}_{2}^{2}+\lambda_{0}\cdot\bm{\beta}^{\top}\hat{\bm{\Gamma}}\bm{\beta}.$ (7)
Eq. (7) shows a connection between featurization and regularization, a perspective explored in several prior works [76, 21, 50, 79]. In particular, $\hat{\bm{\beta}}$ corresponds to the ridgeless limit of generalized ridge regression with generalized penalty matrix $\hat{\bm{\Gamma}}$ [24]. It is not difficult to show that as $\lambda_{0}\to 0$, the predictor $\hat{y}$ can be written as
$\hat{y}(\mathbf{x}_{\mathrm{new}})=\tilde{\mathbf{x}}_{\mathrm{new}}^{\top}\quantity({{\tilde{\mathbf{X}}}}{{\tilde{\mathbf{X}}}})^{+}{{\tilde{\mathbf{X}}}^{\top}}\mathbf{y},$ (8)
where $\tilde{\mathbf{x}}_{\mathrm{new}}:={\mathbf{x}_{\mathrm{new}}^{\top}}\hat{\bm{\Gamma}}^{-1/2}$ and $\tilde{\mathbf{X}}:=\mathbf{X}\hat{\bm{\Gamma}}^{-1/2}$. Using Eq. (5), we may also write $\hat{y}$ as a kernel-ridge predictor with the corresponding kernel $K(\mathbf{x}_{1},\mathbf{x}_{2})=\expectationvalue{\mathbf{x}_{1},\hat{\bm{\Gamma}}^{-1}\mathbf{x}_{2}}$.
In our simple setting, it is intuitively clear that penalizing the ridge regression along a direction (via $\hat{\bm{\Gamma}}$) corresponds to featurizing the data away from that direction (via $\widehat{\mathbf{B}}$). The finding in Remark 2.5 that $r(\cdot)$ is monotonically decreasing shows in addition that this inverse relationship holds along the entire path from no or weak to strong featurization.
@@ 2.3 Related Work
Our model formulation follows a long line of past [66, 65, 60, 41, 62, 15, 70, 20] and concurrent work [8, 33, 72, 77, 36] that assumes that downstream regression tasks share a common linear representation that can be learned upstream. However, most of the existing work focuses on proposing a methodology to learn the ground-truth representation and establish an associated sample complexity bound downstream. Our approach differs from these works in that we characterize the downstream risk exactly. Using this characterization, we study the optimal linear representation which minimizes the average risk over an ensemble of downstream tasks.
Closely related to the studies of asymptotic risk is the effort to reduce the risk by optimizing model parameters. In the context of ridge regression, [48] shows that the double-descent phenomenon can be avoided by optimally tuning the $\ell_{2}$ regularization. The papers [73, 62, 36] study the more general setting of generalized ridge regression. The asymptotic risk is also used for optimal parameter identification beyond ridge regression in [47, 11, 74]. Our work applies this idea to optimal feature construction during a pretraining stage and studies its characteristics from a fine-grained bias-variance tradeoff.

### v2

@@ Abstract
In the transfer learning paradigm models learn useful representations (or features) during a data-rich pretraining stage, and then use the pretrained representation to improve model performance on data-scarce downstream tasks. In this work, we explore transfer learning with the goal of optimizing downstream performance. We introduce a simple linear model that takes as input an arbitrary pretrained feature transform. We derive exact asymptotics of the downstream risk and its fine-grained bias-variance decomposition. We then identify the pretrained representation that optimizes the asymptotic downstream bias and variance averaged over an ensemble of downstream tasks. Our theoretical and empirical analysis uncovers the surprising phenomenon that the optimal featurization is naturally sparse, even in the absence of explicit sparsity-inducing priors or penalties. Additionally, we identify a phase transition where the optimal pretrained representation shifts from hard selection to soft selection of relevant features.
@@ 1 Introduction
Despite the widespread adoption of this paradigm, the mechanisms that produce effective representations remain poorly understood. In particular, one may ask the following question: for a given ensemble of downstream tasks, what constitutes optimal pretrained representation? In this paper, we study a family of downstream linear regression tasks of the form $\mathbf{y}\;=\;\mathbf{X}\,{{\bm{\beta}^{\star}}}+{\bm{\varepsilon}},$ where the true parameter ${{\bm{\beta}^{\star}}}$ factors through a shared representation $\mathbf{B}^{\star}$ via ${{\bm{\beta}^{\star}}}=\mathbf{B}^{\star}{\bm{\alpha}^{\star}}$. We assume that each task-specific coefficient ${\bm{\alpha}^{\star}}$ is drawn from a known prior, and the data matrix $\mathbf{X}$ has Gaussian covariates with covariance $\mathbf{\Sigma}$. Typically, transfer learning proceeds in two steps: (i) obtain an estimator $\tilde{\mathbf{B}}^{\star}$ of the representation $\mathbf{B}^{\star}$ in an upstream, data-rich stage, and (ii) solve the downstream problem $\mathbf{y}=\mathbf{X}\tilde{\mathbf{B}}^{\star}\bm{\alpha}^{\star}+{\bm{\varepsilon}}$ for each new task by estimating $\bm{\alpha}^{\star}$. A key observation of our work is that directly plugging in $\tilde{\mathbf{B}}^{\star}$ can be suboptimal even if $\tilde{\mathbf{B}}^{\star}$ accurately estimates $\mathbf{B}^{\star}$, because $\tilde{\mathbf{B}}^{\star}$ alone does not take into account the covariance structure of $\mathbf{\Sigma}$ or the distribution of $\bm{\alpha}^{\star}$. Instead, one can optimize the feature transform $\widehat{\mathbf{B}}$ to minimize the average downstream risk—the expected risk over the prior of $\bm{\alpha}^{\star}$. Since the downstream data are not available during pretraining, we derive a closed-form expression for the asymptotic downstream risk in the high-dimensional limit, which then becomes a fully differentiable objective for choosing $\widehat{\mathbf{B}}$. We then examine the structure of the optimal $\widehat{\mathbf{B}}$ and its dependence on the shared representation $\mathbf{B}^{\star}$ and the data covariance $\mathbf{\Sigma}$.
@@ 1.1 Contributions
    Deriving the exact asymptotics for the downstream risk and its fine-grained bias-variance decomposition given an arbitrary linear representation from upstream. To elucidate these dependencies, we develop a fine-grained bias-variance decomposition of the population risk, showing how different aspects of $\widehat{\mathbf{B}}$ can reduce bias while inflating variance, and vice versa from a series of ablation studies. In a simpler setting where $\widehat{\mathbf{B}}$ shares eigenvectors with $\mathbf{\Sigma}$, we can explicitly characterize how the balance between aligning to $\mathbf{B}^{\star}$ and leveraging $\mathbf{\Sigma}$ shapes the form of $\widehat{\mathbf{B}}$.
@@
    Proposing optimization methodologies to minimize the asymptotic downstream risk as a function of the pretrained representation; conducting ablation studies comparing the total risk, bias, and variance of the optimally pretrained predictor with predictors with no featurization or ground-truth featurization. We also adapt the methodology for a minimax objective that controls the worst-case performance among downstream tasks.
@@
    Finding that learning both task-relevant features and structures in data covariates are vital, by interpreting the structure of optimal pretraining. Our theoretical and empirical analysis uncovers the surprising phenomenon that the optimal featurization is naturally sparse, even in the absence of explicit sparsity-inducing priors or penalties. Additionally, we uncover a phase transition: when the effective rank of $\mathbf{B}^{\star}$ is below a certain threshold, the optimal transform “hard-selects” principal components (analogous to the classical principal component regression), whereas above that threshold, it “soft-selects” features, smoothly weighting relevant directions for improved performance.
@@
    Conducting numerical experiments when the assumption of shared eigenvectors is lifted. In these more general settings, we empirically observe how the singular vectors of the optimized $\widehat{\mathbf{B}}$ align partially with those of both $\mathbf{B}^{\star}$ and $\mathbf{\Sigma}$.
@@ 2.1 Setting
We model downstream tasks as linear regression problems with Gaussian covariates:
$y\;=\;\mathbf{x}^{\top}\mathbf{\beta}^{\star}+\varepsilon,\quad\quad\mathbf{\beta}^{\star}\;=\;\mathbf{B}^{\star}\bm{\alpha}^{\star},$ (1)
where $\mathbf{x}\sim N(\bm{0},\mathbf{\Sigma})\in\mathbb{R}^{p}$, $\varepsilon\sim N(0,\sigma^{2})$, and $\mathbf{B}^{\star}\in\mathbb{R}^{p\times q}$ is a common, low-rank feature matrix for all downstream tasks. Each task differs by having a specific weight vector $\bm{\alpha}^{\star}\in\mathbb{R}^{q}$ that forms the parameter $\mathbf{\beta}^{\star}=\mathbf{B}^{\star}\bm{\alpha}^{\star}.$ This formulation follows past [62, 61, 56, 39, 58, 14, 66, 19] and concurrent work [8, 30, 68, 73, 34] that assumes that downstream regression tasks share a common linear representation that can be learned upstream.
When specialized downstream data of size $n$ are limited (i.e., $n<p$), directly estimating $\mathbf{\beta}^{\star}$ can suffer high variance. A standard approach leverages the low-rank representation $\mathbf{B}^{\star}$: one can regress on $\bm{\alpha}^{\star}$ via $\mathbf{X}\mathbf{B}^{\star}$, which lives in $\mathbb{R}^{n\times q}$, reducing dimensionality. However, we will see that simply using $\mathbf{B}^{\star}$ to featurize data—even if it were known—does not always minimize downstream risk. The covariance $\mathbf{\Sigma}$ of new data and the prior distribution of $\bm{\alpha}^{\star}$ should also matter.
To capture the variability across different downstream tasks, we impose a prior on $\bm{\alpha}^{\star}$. Specifically, we assume that $\bm{\alpha}^{\star}$ is drawn independently from a distribution $P_{\bm{\alpha}^{\star}}$ with zero mean and covariance $q^{-1}\mathbf{\Sigma}_{\bm{\alpha}^{\star}}$. Our objective is to determine a pretrained feature matrix $\widehat{\mathbf{B}}$ that minimizes the expected downstream risk averaged over this ensemble:
$R^{\mathsf{avg}}=\mathbb{E}_{\bm{\alpha}^{\star}}\Biggl{[}\mathbb{E}_{(\mathbf{X},\mathbf{y},y_{\mathrm{new}},{\mathbf{x}}_{\mathrm{new}})}\Bigl{[}(y_{\mathrm{new}}-\hat{y}_{(\widehat{\mathbf{B}},\mathbf{y},\mathbf{X})}(\mathbf{x}_{\mathrm{new}}))^{2}\Bigr{]}\Biggr{]}.$
Here, the predictor $\hat{y}_{(\widehat{\mathbf{B}},\mathbf{y},\mathbf{X})}$ (introduced in Section 2.2) uses the pretrained feature matrix $\widehat{\mathbf{B}}$ and is fitted on data $(\mathbf{y},\mathbf{X})$ for each downstream task. Averaging the risk over $P_{\bm{\alpha}^{\star}}$ allows us to seek a feature matrix $\widehat{\mathbf{B}}$ that performs well across the entire ensemble of downstream tasks.
To characterize the optimal feature matrix, we will derive the asymptotic limit of the averaged risk, $\mathfrak{R}^{\text{avg}}(\widehat{\mathbf{B}},\mathbf{B}^{\star},\mathbf{\Sigma},\mathbf{\Sigma}_{\bm{\alpha}^{\star}})$ and provide its bias-variance decomposition (see Section 3). We further analyze the optimization problem
$\min_{\widehat{\mathbf{B}}}\mathfrak{R}^{\text{avg}}(\widehat{\mathbf{B}},\mathbf{B}^{\star},\mathbf{\Sigma},\mathbf{\Sigma}_{\bm{\alpha}^{\star}}),$
using both analytical and empirical methods (discussed in Section 5).
Our primary focus in this paper is on deriving and analyzing the optimal $\widehat{\mathbf{B}}$ as a function of $\mathbf{B}^{\star}$, $\mathbf{\Sigma}$. While in practice these quantities may be learned during the upstream pretraining stage, we assume throughout that the ground-truth features $\mathbf{B}^{\star}$ and data covariance $\mathbf{\Sigma}$ are known. The motivation for this assumption is to detach the estimation effects of $\mathbf{B}^{\star}$ and $\mathbf{\Sigma}$ from their role in determining the optimal choice of $\widehat{\mathbf{B}}$. Nevertheless, in Appendix F we consider a simple setting where $\mathbf{B}^{\star}$ is unknown but can be estimated from upstream regression tasks; we then derive the error incurred by using an estimate $\tilde{\mathbf{B}}^{\star}$ in place of $\mathbf{B}^{\star}$ in $\mathfrak{R}^{\text{avg}}$.
@@ 2.2 Ridgeless Regression with Pretrained Representation
In this section, we introduce a linear predictor, $\hat{y}=\hat{y}_{(\widehat{\mathbf{B}},\mathbf{y},\mathbf{X})}$, which is fitted on data $(\mathbf{y},\mathbf{X})$ from a specific downstream task. For our purposes here, the feature matrix $\widehat{\mathbf{B}}\in\mathbb{R}^{p\times k}$ is treated as a fixed, deterministic input. We will delve into the risk decomposition of $\hat{y}$ in Section 3 and, in Section 5, identify the form of $\widehat{\mathbf{B}}$ that optimizes the downstream risk of $\hat{y}$.
@@
For a downstream task with data $(\mathbf{X},\mathbf{y})$, define the empirical loss
@@
$\hat{y}=\mathbf{x}_{\mathrm{new}}^{\top}\hat{\bm{\beta}}\quad\text{where}\quad\quantity(\hat{\bm{\beta}},\hat{\bm{\alpha}})=\lim_{\lambda_{0}\to 0}\operatornamewithlimits{argmin}_{\bm{\beta},\bm{\alpha}}\mathcal{L}(\bm{\beta},\bm{\alpha}).$
Loss function Eq. (2) contains a penalty term $\lambda\|\bm{\beta}-\widehat{\mathbf{B}}\bm{\alpha}\|_{2}^{2}$ that softly enforces the featurization $\hat{\bm{\beta}}=\widehat{\mathbf{B}}\hat{\bm{\alpha}}$. The strength of featurization is controlled by regularization parameters $\lambda_{0},\lambda_{\mathbf{\alpha}},\lambda_{\mathbf{\beta}},\lambda$. One can adjust these parameters to interpolate from strong featurization to no featurization. For instance, we can set $\lambda=0$ which recovers standard ridgeless estimator. More interestingly, we can consider a strong-featurization limit by letting $\lambda_{\mathbf{\beta}}\to 0$ and $\lambda\to+\infty$ for any fixed $\lambda_{\alpha}>0$. It follows from Proposition 2.2 that
@@
Therefore, in this limit, $\hat{y}$ makes predictions by regressing with $\hat{\bm{\alpha}}_{0}$ composed with the featurization generated by $\widehat{\mathbf{B}}$, which is reminiscent of final-layer head-tuning in neural networks [5, 71, 46]. Note that $\hat{\bm{\alpha}}_{0}$ is the minimum $\ell_{2}$-norm solution, and by a well-known result (see e.g. [23]), it is also the limit of gradient flow on the objective $\|\mathbf{y}-\mathbf{X}\widehat{\mathbf{B}}\bm{\alpha}\|^{2}_{2}$ when initialized at zero. These interpretations are the main motivation to consider the ridgeless limit, $\lambda_{0}\to 0$.
Next we present explicit expressions for $\quantity(\hat{\bm{\beta}},\hat{\bm{\alpha}})$. We first write the singular value decomposition of $\widehat{\mathbf{B}}$ as
@@ 2.3 Related Work
Our model formulation follows a long line of past [62, 61, 56, 39, 58, 14, 66, 19] and concurrent work [8, 30, 68, 73, 34] that assumes that downstream regression tasks share a common linear representation that can be learned upstream. However, most of the existing work focuses on proposing a methodology to learn the ground-truth representation and establish an associated sample complexity bound downstream. Additionally, there is often some prior assumptions on low-rank structures or sparsity. Our approach differs from these works in two key aspects (i) we characterize the downstream risk exactly and (ii) we do not impose artificial constraints on the existence of low-dimensional structure or artificially impose sparsity-inducing priors or penalties. Rather, our theoretical and empirical results show that sparsity and feature selection naturally emerges as a consequence of optimizing downstream risk.
The study of asymptotic risk is inherently tied to optimizing model parameters to minimize risk. In ridge regression, [45] demonstrates that careful tuning of $\ell_{2}$ regularization can mitigate the double-descent phenomenon. More relevant to our setting, prior and concurrent works [69, 34] examine generalized ridge regression. [69] analyzes generalized ridge regression in a single-task setting, characterizing risk and classical bias-variance decomposition while studying penalty matrices that optimize bias and variance. However, their asymptotic analysis is limited to cases where the penalty matrix commutes with the data covariance, and their characterization lacks finite-sample error bounds. Moreover, they do not investigate fine-grained bias-variance decomposition, focus on the transfer learning setting, or observe the feature selection and phase transition phenomena we identify. Meanwhile, concurrent work [34] studies generalized ridge regression from a meta-learning perspective, assuming a shared structure among tasks, similar to our work. They characterize asymptotic model risk and analyze the feature matrix that optimizes the asymptotic risk. Their results improve upon [69] by removing the commutativity assumption. However, their setting and analysis differ from ours in two key aspects: (i) we focus on the ridgeless regime, whereas their results require a penalty strength bounded away from zero where there are nontrivial differences in conclusions and analysis. Notably, in their setting, the optimal featurization does not exhibit sparsity (i.e. feature selection); (ii) They do not consider bias-variance decomposition, which is central to our characterization of the feature selection and phase transition phenomena.